# Some remarks on Gorenstein projective precovers

### v2

@@ 4. Generalized Gorenstein flat modules
Note that we have $\mathcal{P}(R)\subseteq\mathcal{FP}_{n}\mbox{-}Flat(R)\subseteq\mathcal{P}(R)^{\wedge}$, thus from [4, Proposition 6.7] the classes $\mathcal{GP}(R)$ and $\mathcal{GP}_{\mathcal{FP}_{n}\mbox{-}Flat(R)}$ are the same. Now from [20, Theorem 3.8] we get that $(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is a complete and hereditary cotorsion pair. ∎
The purpose of this work is to continue with the study of the conditions over a duality pair $(\mathcal{L},\mathcal{A})$ that implies that $\mathcal{GP}_{\mathcal{L}}$ will be special precovering, this conditions will be different to the notion of semidefinable for the class $\mathcal{A}\subseteq\mathrm{Mod}(R^{\mathrm{op}})$. For do this we study interactions between the classes of relative Gorentein flat and relative Gorenstein projective associated to a duality pair $(\mathcal{L},\mathcal{A})$. We begin by recalling some facts and a definition adapted to our setting.
@@
    We see from [12, Corollary 2.20] that for $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{B}\subseteq\mathrm{Mod}(R^{\mathrm{op}})$ and when $\mathcal{GF}_{\mathcal{B}}$ is closed under extensions, then $(\mathcal{GF}_{\mathcal{B}},\mathcal{GF}_{\mathcal{B}}^{\bot_{1}})$ is a hereditary and complete cotorsion pair in $\mathrm{Mod}(R)$. Furthermore from [12, Proposition 3.1] we have that $\mathcal{GF}_{\mathcal{B}}\cap\mathcal{GF}_{\mathcal{B}}^{\bot}=\mathcal{F}(R)\cap\mathcal{F}(R)^{\bot}$.
@@
    Note that when $(\mathcal{L},\mathcal{A})$ is a perfect duality pair, then $\mathcal{P}(R)\subseteq\mathcal{L}$, $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{A}$ and the pair $(\mathcal{P}(R),\mathcal{L})$ is GP-admissible in $\mathrm{Mod}(R)$ (see [4, §3]).
@@
    For a complete duality pair $(\mathcal{L},\mathcal{A})$ we have from [17, Corollary 43] that $(\mathcal{GF}_{\mathcal{A}},\mathcal{GF}_{\mathcal{A}}^{\bot_{1}})$ is a perfect cotorsion pair. In consequence the class $\mathcal{GF}_{\mathcal{A}}$ is closed by extensions and from $\mathrm{(i)}$ and $\mathrm{(ii)}$ also is an hereditary cotorsion pair.
We are ready to prove the third main result, which is a generalization of [13, Main Result].
@@
Let $(\mathcal{X},\mathcal{Y})\subseteq\mathrm{Mod}\times\mathrm{Mod}(R^{\mathrm{op}})$ with the following properties.
@@
    The class $\mathcal{GF}_{\mathcal{Y}}$ is closed under extensions.
Assume that the inclusions $\mathcal{GP}_{\mathcal{X}}\subseteq\mathcal{GF}_{\mathcal{Y}}\subseteq\mathcal{GP}^{\wedge}_{\mathcal{X}}$ are given. Then the class $\mathcal{GP}_{\mathcal{X}}$ is special precovering in $\mathrm{Mod}(R)$.
@@
Take $X\in\mathrm{Mod}(R)$. From Remark 4.6 (i), we have that $(\mathcal{GF}_{\mathcal{Y}},\mathcal{GF}_{\mathcal{Y}}^{\bot})$ is a hereditary and complete cotorsión pair. Thus, there is an exact sequence $0\to Y\to N\to X\to 0$ with $N\in\mathcal{GF}_{\mathcal{Y}}$ and $Y\in\mathcal{GF}_{\mathcal{Y}}^{\bot}$. From the containment $\mathcal{GP}_{\mathcal{X}}\subseteq\mathcal{GF}_{\mathcal{Y}}$, we get $\mathcal{GF}_{\mathcal{Y}}^{\bot}\subseteq\mathcal{GP}_{\mathcal{X}}^{\bot}$. Since $\mathcal{GF}_{\mathcal{Y}}\subseteq\mathcal{GP}^{\wedge}_{\mathcal{X}}$ and $(\mathcal{P}(R),\mathcal{X})$ is a GP-admissible pair, then from [5, Theorem 4.1 (a)] for $N\in\mathcal{GF}_{\mathcal{Y}}$ there is an exact sequence $0\to W\to T\to N\to 0$ with $T\in\mathcal{GP}_{\mathcal{X}}$ and $W\in\mathcal{GP}_{\mathcal{X}}^{\bot}$. We can construct the following p.b digram
@@
and obtain the exact sequence $0\to W\to A\to Y\to 0$, with $W,Y\in\mathcal{GP}_{\mathcal{X}}^{\bot}$, this implies that $A\in\mathcal{GP}_{\mathcal{X}}^{\bot}$. Therefore we obtain the sequence $0\to A\to T\to X\to 0$, with $T\in\mathcal{GP}_{\mathcal{X}}$ and $A\in\mathcal{GP}_{\mathcal{X}}^{\bot}$. This is, $T\to X$ is a special $\mathcal{GP}_{\mathcal{X}}$-precover. ∎
@@
Let $R$ be a left n-perfect ring and such that $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R^{\mathrm{op}})^{\wedge}$. Then, the class of Ding projective $R$-modules $\mathcal{DP}(R)$ is special precovering in $\mathrm{Mod}(R)$.
@@
We will verify the conditions of the Theorem 4.7 with the pair $(\mathcal{F}(R),\mathcal{I}(R^{\mathrm{op}}))$. To this end note that $(\mathcal{P}(R),\mathcal{F}(R))$ is a GP-admissible pair and $\mathcal{GP}_{\mathcal{F}}$ is precisely the class of Ding-projective $R$-modules $\mathcal{DP}(R)$, while $\mathcal{GF}_{\mathcal{I}(R^{\mathrm{op}})}$ is the usual class of Gorenstein flat $R$-modules $\mathcal{GF}(R)$. From [27, Corollary 2.6] we have the containment $\mathcal{GP}(R)\subseteq\mathcal{GF}(R)$ and always is true that $\mathcal{DP}(R)\subseteq\mathcal{GP}(R)$, therefore $\mathcal{DP}(R)\subseteq\mathcal{GF}(R)$. Also, from [9, Proposition 3.9], we have that $\mathcal{GF}(R)\subseteq\mathcal{P}\mathcal{GF}(R)^{\wedge}_{n}$ and by [17, Corollary 14] the containment $\mathcal{P}\mathcal{GF}(R)\subseteq\mathcal{DP}(R)$ is always true. All this give us $\mathcal{GF}(R)\subseteq\mathcal{P}\mathcal{GF}(R)^{\wedge}\subseteq\mathcal{DP}(R)^{\wedge}$. ∎
Until now, from Proposition 4.2 we know that when $(\mathcal{L},\mathcal{A})$ is a symmetric duality the containment $\mathcal{GP}_{\mathcal{L}}\subseteq\mathcal{GF}_{\mathcal{A}}$ is given. By Remark 4.6 if $(\mathcal{L},\mathcal{A})$ is perfect, then are fulfilled the conditions $\mathrm{(i),(ii)}$ in the previous Theorem 4.7. While if $(\mathcal{L},\mathcal{A})$ is complete then $\mathrm{(iii)}$ in Theorem 4.7 is also true. In the present generality, we have the following result.
Proposition 4.9.
Let $(\mathcal{L},\mathcal{A})$ be a complete duality pair in $\mathrm{Mod}(R)$ and $R$ a left n-perfect ring. Then the class $\mathcal{GP}_{\mathcal{L}}$ is special precovering on $\mathrm{Mod}(R)$.
@@
Since $(\mathcal{L},\mathcal{A})$ complete, also is symmetric and perfect. Thus, the conditions in Theorem 4.7 are fulfilled, except for the containment $\mathcal{GF}_{\mathcal{A}}\subseteq\mathcal{GP}^{\wedge}_{\mathcal{L}}$. To prove such a containment we must assume that $R$ is $n$-perfect, that is $\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}_{n}$.
@@
with $Z_{j}(\mathbf{P}_{i})\in\mathcal{P}(R)$ for each $j$ and where $Z_{j}(\mathbf{C})\in\mathcal{P}\mathcal{GF}_{\mathcal{A}}=\mathcal{GP}_{\mathcal{L}}$ by Proposition 4.2. Thus for $j=0$ we get the desired $\mathcal{GP}_{\mathcal{L}}$-resolution. ∎
@@
Note that from the proof of Proposition 4.9 we can see that when $(\mathcal{L},\mathcal{A})$ is a symmetric duality pair and $R$ is $n$-perfect, then $\mathcal{GF}_{\mathcal{A}}\subseteq{[\mathcal{GP}_{\mathcal{L}}]}^{\wedge}_{n}$.
When the ring $R$ is right coherent then the duality pair $(\mathcal{F}(R),\mathcal{I}(R^{\mathrm{op}}))$ es is symmetric, and since $(\mathcal{F},\mathcal{F}^{\bot})$ is always a complete cotorsion pair it follows that $(\mathcal{F}(R),\mathcal{I}(R^{\mathrm{op}}))$ is a complete duality pair. Thus, when $R$ is also left $n$-perfect we recover the main result in [13], as follows .
Proposition 4.11.
Let $R$ be a ring right coherent and left $n$-perfect. The following statements are true.
  • (a)
    The class of Ding projective $R$-modules $\mathcal{DP}(R)$ is special precovering in $\mathrm{Mod}(R)$
  • (b)
    The class of Gorenstein projective $R$-modules $\mathcal{GP}(R)$ is special precovering in $\mathrm{Mod}(R)$.
Proof.
(a) Since $R$ is right coherent, the pair $(\mathcal{F}(R),\mathcal{I}(R^{\mathrm{op}}))$ is a complete duality pair, then from Proposition 4.9 we get that $\mathcal{GP}_{\mathcal{F}}$ is special precovering, but this class is precisely the class of Ding projectives.
(b) Note that we have $\mathcal{P}(R)\subseteq\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}_{n}$, thus from [4, Proposition 6.7] we get $\mathcal{GP}_{\mathcal{F}}=\mathcal{DP}(R)$, therefore the result it follows from (a). ∎
Since from [17, Proposition 23] the condition of being special precovering is sufficient to obtain a complete and hereditary cotorsion pair, we have the following result, which is a summary of what we have proved so far.
@@
    If for some $n\geq 2$ the containment $\mathcal{FP}_{n}\mbox{-}Flat(R)\subseteq\mathcal{P}(R)^{\wedge}$ is given, then $(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is a complete and hereditary cotorsion pair in $\mathrm{Mod}(R)$.
@@
    If $R$ is left n-perfect and $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R^{\mathrm{op}})^{\wedge}$, then $(\mathcal{DP}(R),\mathcal{DP}^{\bot})=(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is a complete and hereditary cotorsion pair in $\mathrm{Mod}(R)$.
@@
    If $(\mathcal{L},\mathcal{A})$ is a complete duality pair in $\mathrm{Mod}(R)$ and $R$ a left n-perfect, then the pair $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot})$ is a complete and hereditary cotorsion pair in $\mathrm{Mod}(R)$.
Now we will see that it is possible to get more cotorsion pairs induced by these. In fact in each case of the above proposition we will obtain a family of cotorsion pairs. In order to obtain such cotorsion pairs we need the following lemma.
@@
From [10, Theorem 7.4.6], for $H$ there is an exact sequence $0\to K^{\prime}\to H^{\prime}\to H\to 0$, with $H^{\prime}\in\mathcal{P}(R)^{\wedge}_{m}$ and $K^{\prime}\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}$. Note that $H,H^{\prime}\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ (since $\mathcal{P}(R)\subseteq\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ implies $\mathcal{P}(R)^{\wedge}_{m}\subseteq\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ by the dual of [5, Lemma 2.6]). This implies that $K^{\prime}\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$, i.e. $K^{\prime}\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$, so that $\mathrm{Ext}^{1}_{R}(H,K^{\prime})=0$. That is $0\to K^{\prime}\to H^{\prime}\to H\to 0$ splits, therefore $H\in\mathcal{P}(R)^{\wedge}_{m}$. Thus, the exact sequence $0\to M\to H\to Q\to 0$ fulfils with the conditions of [5, Corollary 4.3 (c)]. ∎
@@
Form Lemma 4.13 we know that
$[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$ =([$={{}^{\bot_{1}}\displaystyle{\left([\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}\right)}}$ and $[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ ([$\subseteq{{}^{\bot_{1}}([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m})}.$
We have for the other hand that $\mathcal{P}(R)^{\wedge}_{m}\cup\mathcal{GP}_{\mathcal{L}}\subseteq[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}\cup\mathcal{GP}_{\mathcal{L}}=[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$. Then, taking orthogonals
@@
This implies that $\left([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}\right)^{\bot_{1}}=(\mathcal{P}(R)^{\wedge}_{m})^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. Since the class $[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$ is closed by kernels of epimorphisms [5, Corollary 4.10], we conclude that such cotorsion pair is hereditary. It remains to prove that such a pair is complete. We know that $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ is complete as well as $(\mathcal{P}(R)^{\wedge}_{m},[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}})$ [10, Theorem 7.4.6], thus for $M\in\mathrm{Mod}(R)$ there is an exact sequence $0\to K\to H\to M\to 0\to 0$ with $H\in\mathcal{GP}_{\mathcal{L}}$ and $K\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. And to this $K$ there is $0\to K\to S\to E\to 0$ with $S\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}$ and $E\in\mathcal{P}(R)^{\wedge}_{m}$. Consider the following p.o. diagram
@@
As a natural consequence we have a Hovey triple in $\mathrm{Mod}(R)$.
@@
Whose homotopy category is the stable category $[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}\cap[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}/([\mathcal{P}(R)^{\wedge}_{m}]\cap[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}})$.

### v3

@@ 3.1. Precovers from generalized Gorenstein $R$-modules
Note that we have $\mathcal{P}(R)\subseteq\mathcal{FP}_{n}\mbox{-}Flat(R)\subseteq\mathcal{P}(R)^{\wedge}$, thus from [4, Proposition 6.7] the classes $\mathcal{GP}(R)$ and $\mathcal{GP}_{\mathcal{FP}_{n}\mbox{-}Flat(R)}$ are the same. Now from [28, Theorem 3.8] we get that $(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is a complete and hereditary cotorsion pair. ∎
Based on the above result we are interested in the study of the conditions over a duality pair $(\mathcal{L},\mathcal{A})$, that implies that $\mathcal{GP}_{\mathcal{L}}$ will be special precovering. This conditions will be different to the notion of semidefinable for the class $\mathcal{A}\subseteq\mathrm{Mod}(R^{\mathrm{op}})$. For do this, we study interactions between the classes of relative Gorenstein flat and relative Gorenstein projective associated to a duality pair $(\mathcal{L},\mathcal{A})$. We begin by recalling some facts and a definition adapted to our setting.
@@
    We see from [18, Corollary 2.20] that for a class of right $R$-modules $\mathcal{B}$ such that $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{B}$ which satisfies that $\mathcal{GF}_{\mathcal{B}}$ is closed under extensions, then the pair $(\mathcal{GF}_{\mathcal{B}},\mathcal{GF}_{\mathcal{B}}^{\bot_{1}})$ is a hereditary and complete cotorsion pair.
@@
    Note that when $(\mathcal{L},\mathcal{A})$ is a perfect duality pair, then the containments $\mathcal{P}(R)\subseteq\mathcal{L}$, and $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{A}$ are given. Furthermore the pair $(\mathcal{P}(R),\mathcal{L})$ is GP-admissible (see [4, §3]).
@@
    For a complete duality pair $(\mathcal{L},\mathcal{A})$ we know from [23, Corollary 43] that $(\mathcal{GF}_{\mathcal{A}},\mathcal{GF}_{\mathcal{A}}^{\bot_{1}})$ is a perfect cotorsion pair. In consequence the class $\mathcal{GF}_{\mathcal{A}}$ is closed by extensions and by $\mathrm{(i)}$ and $\mathrm{(ii)}$ also is an hereditary cotorsion pair.
We are ready to prove a useful result, which is a generalization of [19, Main Result].
@@
Let $(\mathcal{X},\mathcal{Y})\subseteq\mathrm{Mod}\times\mathrm{Mod}(R^{\mathrm{op}})$ be with the following properties.
@@
    The class $\mathcal{GF}_{\mathcal{Y}}$ is closed by extensions.
Assume that the inclusions $\mathcal{GP}_{\mathcal{X}}\subseteq\mathcal{GF}_{\mathcal{Y}}\subseteq\mathcal{GP}^{\wedge}_{\mathcal{X}}$ are given. Then the class $\mathcal{GP}_{\mathcal{X}}$ is special precovering.
@@
Take $X\in\mathrm{Mod}(R)$. From Facts 3.9 (i), we have that $(\mathcal{GF}_{\mathcal{Y}},\mathcal{GF}_{\mathcal{Y}}^{\bot})$ is a hereditary and complete cotorsión pair. Thus, there is an exact sequence $0\to Y\to N\to X\to 0$ with $N\in\mathcal{GF}_{\mathcal{Y}}$ and $Y\in\mathcal{GF}_{\mathcal{Y}}^{\bot}$. From the containment $\mathcal{GP}_{\mathcal{X}}\subseteq\mathcal{GF}_{\mathcal{Y}}$, we get $\mathcal{GF}_{\mathcal{Y}}^{\bot}\subseteq\mathcal{GP}_{\mathcal{X}}^{\bot}$. Since $\mathcal{GF}_{\mathcal{Y}}\subseteq\mathcal{GP}^{\wedge}_{\mathcal{X}}$ and $(\mathcal{P}(R),\mathcal{X})$ is a GP-admissible pair, then from [5, Theorem 4.1 (a)] for $N\in\mathcal{GF}_{\mathcal{Y}}$ there is an exact sequence $0\to W\to T\to N\to 0$ with $T\in\mathcal{GP}_{\mathcal{X}}$ and $W\in\mathcal{GP}_{\mathcal{X}}^{\bot}$. We can construct the following p.b digram
@@
and obtain the exact sequence $0\to W\to A\to Y\to 0$, with $W,Y\in\mathcal{GP}_{\mathcal{X}}^{\bot}$, this implies that $A\in\mathcal{GP}_{\mathcal{X}}^{\bot}$. Therefore we obtain the exact sequence $0\to A\to T\to X\to 0$, with $T\in\mathcal{GP}_{\mathcal{X}}$ and $A\in\mathcal{GP}_{\mathcal{X}}^{\bot}$. That is, $T\to X$ is a special $\mathcal{GP}_{\mathcal{X}}$-precover. ∎
As application we have the following.
@@
Let $R$ be a ring such that $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R^{\mathrm{op}})^{\wedge}$ and $\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}$. Then, the class of Ding projective $R$-modules $\mathcal{DP}(R)$ is special precovering.
@@
We will verify the conditions of Theorem 3.10 with the pair $(\mathcal{F}(R),\mathcal{I}(R^{\mathrm{op}}))$. To this end note that $(\mathcal{P}(R),\mathcal{F}(R))$ is a GP-admissible pair and $\mathcal{GP}_{\mathcal{F}}$ is precisely the class of Ding-projective $R$-modules $\mathcal{DP}(R)$, while $\mathcal{GF}_{\mathcal{I}(R^{\mathrm{op}})}$ is the usual class of Gorenstein flat $R$-modules $\mathcal{GF}(R)$. From [35, Corollary 2.6] we have the containment $\mathcal{GP}(R)\subseteq\mathcal{GF}(R)$ and always is true that $\mathcal{DP}(R)\subseteq\mathcal{GP}(R)$, therefore $\mathcal{DP}(R)\subseteq\mathcal{GF}(R)$. Also, from [11, Proposition 3.9], we have that $\mathcal{GF}(R)\subseteq\mathcal{P}\mathcal{GF}(R)^{\wedge}$ and by [23, Corollary 14] the containment $\mathcal{P}\mathcal{GF}(R)\subseteq\mathcal{DP}(R)$ is always true. All this give us $\mathcal{GF}(R)\subseteq\mathcal{P}\mathcal{GF}(R)^{\wedge}\subseteq\mathcal{DP}(R)^{\wedge}$. ∎
Is know from [12, Theorem 3.8] and [20, Corollaries 4.5 and 4.6] that when $R$ is a Ding-Chen ring the class $\mathcal{DP}(R)$ of Ding projective $R$-modules is a class special precovering, the result above and Proposition 3.3 exhibits other conditions of when this occurs and shows the usefulness of Theorem 3.10.
Until now, from Proposition 3.5 we know that when $(\mathcal{L},\mathcal{A})$ is a symmetric duality the containment $\mathcal{GP}_{\mathcal{L}}\subseteq\mathcal{GF}_{\mathcal{A}}$ is given. By Facts 3.9, if $(\mathcal{L},\mathcal{A})$ is perfect, then are fulfilled the conditions $\mathrm{(i),(ii)}$ in the previous Theorem 3.10. While if $(\mathcal{L},\mathcal{A})$ is complete then $\mathrm{(iii)}$ in Theorem 3.10 is also true. In the present generality, we declare the following result.
Theorem 3.12.
Let $(\mathcal{L},\mathcal{A})$ be a complete duality pair in $\mathrm{Mod}(R)$ and $R$ a left n-perfect ring. Then, the class $\mathcal{GP}_{\mathcal{L}}$ is special precovering.
@@
Since $(\mathcal{L},\mathcal{A})$ is complete, also is symmetric and perfect. Thus, the conditions in Theorem 3.10 are fulfilled, except for the containment $\mathcal{GF}_{\mathcal{A}}\subseteq\mathcal{GP}^{\wedge}_{\mathcal{L}}$. To prove such a containment we must assume that $R$ is $n$-perfect, that is $\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}_{n}$.
@@
with $Z_{j}(\mathbf{P}_{i})\in\mathcal{P}(R)$ for each $j$ and where $Z_{j}(\mathbf{C})\in\mathcal{P}\mathcal{GF}_{\mathcal{A}}=\mathcal{GP}_{\mathcal{L}}$ by Proposition 3.5. Thus, for $j=0$ we get the desired $\mathcal{GP}_{\mathcal{L}}$-resolution. ∎
@@
Note that, from the proof of Theorem 3.12 we see that when $(\mathcal{L},\mathcal{A})$ is a symmetric duality pair and $R$ is $n$-perfect, then $\mathcal{GF}_{\mathcal{A}}\subseteq{[\mathcal{GP}_{\mathcal{L}}]}^{\wedge}_{n}$.
When $R$ is a right coherent ring the duality pair $(\mathcal{F}(R),\mathcal{I}(R^{\mathrm{op}}))$ es is symmetric, and since $(\mathcal{F},\mathcal{F}^{\bot})$ is always a complete cotorsion pair, it follows that $(\mathcal{F}(R),\mathcal{I}(R^{\mathrm{op}}))$ is a complete duality pair. Thus, in the case when $R$ is also left $n$-perfect we recover from Theorem 3.12 the main result in [19].
Since the condition of being special precovering is enough to obtain a complete and hereditary cotorsion pair [23, Proposition 23], we give in the following result a summary in such terms of what we have proved so far.
@@
    If the containment $\mathcal{GP}(R)\subseteq\mathcal{GF}(R)$ is given, then $(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is a complete and hereditary cotorsion pair.
@@
    For $R$ a right coherent ring, the pair $(\mathcal{DP}(R),\mathcal{DP}(R)^{\bot})$ is a complete and hereditary cotorsion pair in $\mathrm{Mod}(R)$. Furthermore if $\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}$ then $(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is also an complete and hereditary cotorsion pair.
@@
    If for some $n\geq 2$ the containment $\mathcal{FP}_{n}\mbox{-}Flat(R)\subseteq\mathcal{P}(R)^{\wedge}$ is given, then $(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is a complete and hereditary cotorsion pair.
  • (iv)
    If $R$ is left n-perfect and $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R^{\mathrm{op}})^{\wedge}$, then $(\mathcal{DP}(R),\mathcal{DP}^{\bot})=(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is a complete and hereditary cotorsion pair.
  • (v)
    If $(\mathcal{L},\mathcal{A})$ is a complete duality pair in $\mathrm{Mod}(R)$ and $R$ a left n-perfect, then the pair $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot})$ is a complete and hereditary cotorsion pair.
Proof.
(i) It follows from Remark 3.1, or equivalently by Proposition One-Main.
(ii) This is Proposition 3.3.
(iii) Comes from Proposition 3.7.
(iv) This is Corollary 3.11, where the equality is given from [5, Proposition 6.7].
(v) It follows from Theorem 3.12. ∎
An interesting phenomena comes from the existence of the cotorsion pair $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot})$ since it is possible to get more cotorsion pairs induced by these. In fact, in each case of the proposition above, we will obtain a family of cotorsion pairs. In order to obtain such cotorsion pairs we need the following lemma.
@@
From [15, Theorem 7.4.6], for $H$ there is an exact sequence $0\to K^{\prime}\to H^{\prime}\to H\to 0$, with $H^{\prime}\in\mathcal{P}(R)^{\wedge}_{m}$ and $K^{\prime}\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}$. Note that $H,H^{\prime}\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ (since $\mathcal{P}(R)\subseteq\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ implies $\mathcal{P}(R)^{\wedge}_{m}\subseteq\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ by the dual of [5, Lemma 2.6]). This implies that $K^{\prime}\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$, i.e. $K^{\prime}\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$, so that $\mathrm{Ext}^{1}_{R}(H,K^{\prime})=0$. That is $0\to K^{\prime}\to H^{\prime}\to H\to 0$ splits, therefore $H\in\mathcal{P}(R)^{\wedge}_{m}$. Thus, the exact sequence $0\to M\to H\to Q\to 0$ fulfils with the conditions of [5, Corollary 4.3 (c)]. ∎
@@
From Lemma 3.15 we know that following;
$[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$ $={{}^{\bot_{1}}\displaystyle{\left([\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}\right)}}$ and $[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$ $\subseteq{([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m})^{\bot_{1}}}.$
For the other hand we have that $\mathcal{P}(R)^{\wedge}_{m}\cup\mathcal{GP}_{\mathcal{L}}\subseteq[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}\cup\mathcal{GP}_{\mathcal{L}}=[\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$. Then, taking orthogonals
@@
This implies that $\left([\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}\right)^{\bot_{1}}=(\mathcal{P}(R)^{\wedge}_{m})^{\bot_{1}}\cap\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. Since the class $\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}$ is closed by kernels of epimorphisms [5, Corollary 4.10], we conclude that such cotorsion pair is hereditary. It remains to prove that such a pair is complete. We know that $(\mathcal{GP}_{\mathcal{L}},\mathcal{GP}_{\mathcal{L}}^{\bot_{1}})$ is complete as well as $(\mathcal{P}(R)^{\wedge}_{m},[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}})$ [15, Theorem 7.4.6], thus for $M\in\mathrm{Mod}(R)$ there is an exact sequence $0\to K\to H\to M\to 0\to 0$ with $H\in\mathcal{GP}_{\mathcal{L}}$ and $K\in\mathcal{GP}_{\mathcal{L}}^{\bot_{1}}$. And to this $K$ there is $0\to K\to S\to E\to 0$ with $S\in[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}$ and $E\in\mathcal{P}(R)^{\wedge}_{m}$. Consider the following p.o. diagram
@@
Thus, in each of the conditions of Proposition 3.14, we get a family of cotorsion pairs. As an expected consequence we also obtain a Hovey triple in $\mathrm{Mod}(R)$.
@@
Whose homotopy category is the stable category $\mathcal{GP}_{\mathcal{L}}]^{\wedge}_{m}\cap[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}}/([\mathcal{P}(R)^{\wedge}_{m}]\cap[\mathcal{P}(R)^{\wedge}_{m}]^{\bot_{1}})$.